# A Law of Next-Token Prediction in Large Language Models

### v1

@@ Acknowledgments
Fig. S12: Pre-training data can affect how the law of equi-learning behaves. The same GPT-2 models pre-trained on C4 and OSCAR are probed with the OpenWebText dataset.
Pre-training data analysis.
As illustrated in Fig. S12, GPT-2 models pre-trained on two different datasets, C4 and OSCAR, exhibit distinct behaviors regarding the emergence of the law of equi-learning when evaluated on the same probing dataset (i.e., OpenWebText). Specifically, models pre-trained with OSCAR demonstrate more noise in the law’s emergence compared to those pre-trained with C4. This discrepancy is likely attributable to the higher noise levels in OSCAR, stemming from its less stringent deduplication. These models were released by [26]. Our findings underscore the significant impact of pre-training data quality on the emergence of the law of equi-learning, suggesting that higher quality pre-training data may result in a more pronounced manifestation of this law.

### v3

@@ Acknowledgments
Fig. S12: Pre-training data can affect how the law of equi-learning behaves. The same GPT-2 models pre-trained on C4 and OSCAR are probed with the OpenWebText dataset. The x-axis denotes the layer index, while the y-axis (log scale) shows the prediction residual (PR) as defined in Eq. 2.
Pre-training data analysis. As illustrated in Fig. S12, GPT-2 models pre-trained on two different datasets, C4 and OSCAR, exhibit distinct behaviors regarding the emergence of the law of equi-learning when evaluated on the same probing dataset (i.e., OpenWebText). Specifically, models pre-trained with OSCAR demonstrate more noise in the law’s emergence compared to those pre-trained with C4. This discrepancy is likely attributable to the higher noise levels in OSCAR, stemming from its less stringent deduplication. These models were released by [28]. Our findings underscore the significant impact of pre-training data quality on the emergence of the law of equi-learning, suggesting that higher quality pre-training data may result in a more pronounced manifestation of this law.